# Multi-scale Latent Point Consistency Models for 3D Shape Generation

### v1

@@ 4 Experiments
Figure 3: Unconditional shape generation with 2,048 points for the classes airplane, car, and chair using class-specific models, trained on the ShapeNet dataset from PointFlow with global normalization.
Figure 4: Qualitative visualizations of high-fidelity and diverse 3D point cloud generation.
4.1 Datasets & Metrics
Datasets. To compare MLPCM against existing methods, we use ShapeNet [4], the most widely used dataset to benchmark 3D shape generative models. Following previous works [56, 55, 61, 30], we train on three categories: airplane, chair, car and primarily rely on PointFlow’s [55] dataset splits and preprocssing, which normalizes the data globally across the whole dataset. However, some baselines require per-shape normalization [19, 58, 3, 10]; hence, we also train on such data. In addition, following [56], we also perform the task of multi-class unconditional 3D shape generation on the ShapeNet-Vol [38, 37] dataset, where the data is normalized per-shape.
@@ 4.1 Datasets & Metrics
Evaluation. Following recent works [55, 61, 56], we use 1-NNA, with both Chamfer distance (CD) and earth mover distance (EMD), as our main metric. It quantifies the similarity between the distribution of generated shapes and that of shapes from the validation set and measures both quality and diversity [55].
4.2 Single-Class 3D Point Cloud Generation
@@
Implementation Our implementation is based on the PyTorch [36]. The input point cloud size is $2048\times 3$. Both our VAE encoder and decoder are built upon PVCNN [26], consisting of 4 layers of SA modules and FP modules. The voxel grid sizes of PVC at different scales are $32$, $16$, $8$, and $8$, respectively. We adopt the Farthest Point Sampling (FPS) algorithm for sampling in the Grouper block, with the sampled center points being $1024$, $256$, $64$, and $16$. KNN (K-Nearest Neighbors) is used to aggregate local neighborhood features, with $32$ neighbors for each point in the Grouper. We apply a dropout rate of $0.1$ to all dropout layers in the VAE. Following [56], we initialize our VAE model so that it acts as an identity mapping between the input, latent space, and reconstruction points at the start of training. We achieve this by reducing the variance of the encoder and correspondingly weighting the skip connections. Our latent prior also consists of 4 layers of SA modules and FP modules. The time embedding and multi-scale latent representations are concatenated with the point features at the input of each corresponding SA and FP layer. We adopt the hybrid denoising score network parameterization following [56].
@@
Results We use the Adam optimizer [17] to train our model, where the VAE is trained for $8,000$ epochs and the diffusion prior for $24,000$ epochs, with a batch size of $128$. We set $T=1,000$ as the time steps of the diffusion process. For latent consistency distillation, we train MLPCM for $100$K iterations with a batch size of $128$, a learning rate of $2\times 10^{-4}$, and an EMA rate $\mu=0.99995$. We validate our model’s results under two settings: normalizing data across the entire dataset and normalizing each shape individually. Samples from MLPCM are shown in Figure 3 and Figure  4, and quantitative results are provided in Table 1 and Table 2. Note that TM stands for Teacher Model and LCM stands for Latent Consistency Models. MLPCM outperforms all baselines and achieves state-of-the-art performance across all categories and dataset versions. Compared to key baselines like DPM, PVD, and LION, our samples are diverse and have a better visual quality.
@@
Table 1: Generation metrics (1-NNA↓) on airplane, chair, car categories from ShapeNet dataset from PointFlow [55]. Training and test data normalized globally into [-1, 1]. TM stands for Teacher Model and LCM stands for Latent Consistency Models.
@@
Table 2: Generation results (1-NNA↓) on ShapeNet dataset from PointFlow [55]. All data normalized individually into [-1, 1].
@@
4.3 Multi-Class 3D Shape Generation
We validated the model’s generation performance on the Shapenet-Vol dataset, where each shape is normalized, meaning the point coordinates are bounded within $[-1,1]$. We report the quantitative results in Table 3, comparing our method with various strong baselines under the same setting. We found that the proposed multi-scale latent diffusion model significantly outperforms all baselines, and MLPCM greatly improves sampling efficiency while maintaining performance.
Table 3: Results of joint training on 13 classes of ShapeNet-vol.
@@ 4.3 Multi-Class 3D Shape Generation
4.4 Sampling Time
The proposed multi-scale latent diffusion model synthesizes shapes using 1000 steps of DDPM, while MLPCM synthesizes them using only 1-4 steps, generating high-quality shapes in under 0.5 seconds. This makes real-time interactive applications feasible. In Table 4, we compare the sampling time and quality of generating a point cloud sample (2048 points) from noise using different DDPM, DDIM [46] sampling methods, and MLPCM. When using $\leq 10$ steps, DDIM’s performance degrades significantly, whereas MLPCM can produce visually high-quality shapes in just a few steps or even one step.
Table 4: The sampling rates and generation quality of our model with various methods. Our approach achieves high-quality shape generation within 0.5 seconds while maintaining quality.
@@ 4.4 Sampling Time
4.5 Ablation Study
Ablation on the Key Design Components. We perform the ablation study on the Car category. For the configuration without multi-scale latent representations, we consider two scenarios: one using only point-level latent features to capture fine-grained details and another using only shape-level features to focus on global structure. This allows us to assess the impact of each representation scale individually. The model sizes are kept nearly identical across all settings to ensure fair comparison. As shown in Table 5, the full model, integrating multi-scale latent representations and the 3D spatial attention mechanism, consistently outperforms the simplified settings across all evaluation metrics. This demonstrates the effectiveness of combining fine-grained local information with global shape context, along with the attention mechanism, in achieving more accurate and coherent reconstructions.
Table 5: Ablation studies on key 3D design components on the car category. PL and SL represent the point-level latents and shape-level latents, respectively, and 3DSA represents the 3D Spatial Attention Mechanism.
@@ 4.5 Ablation Study
Table 6: Ablation on the number of extra dimension $D_{l}$ of the latent points, on the car category.
@@
Ablation on Extra Dimension for Latent Points. We ablate the additional latent point dimension $D_{l}$ on the Car category in Table 6. We experimented with several different additional dimensions for the latent points, ranging from 0 to 5, where $D_{l}=1$ provided the best overall performance. As the additional dimensions increase, we observe a decrease in the 1-NNA score. We thus use $D_{l}=1$ for all other experiments.
5 Conclusion

### v2

@@ IV Experiments
Figure 3: Unconditional shape generation with 2,048 points for the classes airplane, car, and chair using class-specific models, trained on the ShapeNet dataset from PointFlow with global normalization. Our samples better represent component relationships, such as the seat-leg connection and the wing-fuselage connection, while competing approaches exhibit smaller gaps, fused components, or asymmetries between mirrored components.
Figure 4: Qualitative visualizations of high-fidelity and diverse 3D point cloud generation. Generated shapes preserve physically plausible structure and global topology while varying fine details, demonstrating diversity without sacrificing symmetry.
Figure 5: Comparison between generated shapes and retrieved nearest neighbors (NN) in the training dataset. The generated shapes closely match the ground-truth shapes in overall topology, yet differ in fine-grained details, demonstrating diversity.
Figure 6: Voxel-conditioned synthesis for interactive generation. Given a coarse voxel input, our model produces diverse 3D shapes that adhere to the occupied regions while refining geometry into clean surfaces and coherent parts.
IV-A Datasets
For the unconditional 3D point cloud generation task, we utilize the ShapeNet dataset [70] as pre-processed and provided by PointFlow [2], ensuring consistency with prior studies. To facilitate fair comparisons with existing baseline methods [2, 71, 3, 17, 1], we focus on three commonly used categories: airplane, chair, and car. Each shape in the dataset contains 15,000 points, and during training, we randomly sample 2,048 points per shape in each iteration. The training dataset includes 2,832 shapes for airplanes, 4,612 for chairs, and 2,458 for cars.
We evaluate sample quality using a standard reference set, which consists of 405, 662, and 352 shapes for airplanes, chairs, and cars, respectively. Consistent with the normalization approach in PointFlow [2] and LION [3], the data is normalized globally across all shapes. Specifically, we calculate the mean values for each axis and a single standard deviation across all axes based on the entire training set. This ensures that the input data is standardized for training and comparison purposes.
While reproducing baseline methods on the ShapeNet dataset provided by PointFlow [2], we observed that certain approaches [3, 72, 73, 65, 74] utilize per-shape normalization. This involves calculating the mean for each axis individually for each shape and determining the scale based on the maximum extent across all axes for that shape. Consequently, the point coordinates are constrained within the range of [-1, 1]. For comparisons with these methods, we adhere to the same normalization strategy during both training and evaluation of our model. It is worth noting that these varying normalization schemes lead to fundamentally different generative modeling tasks. Therefore, ensuring clarity about these setups is crucial for fair and meaningful comparisons.
IV-B Implementation Details
Training Details Our teacher model is implemented in PyTorch [75]. The VAE is trained for around 8000 epochs with an initial learning rate of 0.001, and we use Adam [76] to optimize the entire model. The batch size is set to 128, the latent point dimension is 4, and the skip weight is set to 0.01. The diffusion prior is trained for approximately 24,000 epochs with an initial learning rate of 0.0002, using 1000 time steps for the diffusion process, with a linear schedule for $\beta_{t}$. We train MLPCM for 100100K iterations with a batch size of $128$, a learning rate of $2\times 10^{-4}$, and an EMA rate $\mu=0.99995$. Both the VAE and the diffusion prior are trained on 8 NVIDIA A40 GPUs.
For the mixed denoising score network parametrization used in our teacher model, we follow the approach of [77]. Specifically, the denoising model is parameterized as an analytically optimal denoising network for a normal data distribution, augmented with a neural network-based correction. This design is particularly advantageous when the modeled distribution is close to normal, as is the case here. In the first training stage, all latent encodings are regularized to approximate a standard normal distribution due to the Kullback-Leibler regularization in the VAE objective. Our implementation of the mixed denoising score network strictly adheres to the method described by [77], and we direct readers to their work for additional details.
For the latent consistency model ,to align with the discrete-time schedule in our multi-scale latent diffusion model, which synthesizes shapes over 1000 DDPM steps, we employ DDIM [31] as the PF-ODE solver.
@@ IV-B Implementation Details
Evaluation Metrics Following recent works [2, 71, 3], we use 1-NNA, with both Chamfer distance (CD) and earth mover distance (EMD), as our main metric. It quantifies the similarity between the distribution of generated shapes and that of shapes from the validation set and measures both quality and diversity [2].
IV-C Single-Class 3D Point Cloud Generation
@@
Implementation Our implementation is based on the PyTorch. The input point cloud size is $2048\times 3$. Both our VAE encoder and decoder are built upon PVCNN [68], consisting of 4 layers of SA modules and FP modules. The voxel grid sizes of PVC at different scales are $32$, $16$, 88, and $8$, respectively. We adopt the Farthest Point Sampling (FPS) algorithm for sampling in the Grouper block, with the sampled center points being $1024$, $256$, $64$, and $16$. KNN (K-Nearest Neighbors) is used to aggregate local neighborhood features, with $32$ neighbors for each point in the Grouper. We apply a dropout rate of $0.1$ to all dropout layers in the VAE. Following [3], we initialize our VAE model so that it acts as an identity mapping between the input, latent space, and reconstruction points at the start of training. We achieve this by reducing the variance of the encoder and correspondingly weighting the skip connections. Our latent prior also consists of 4 layers of SA modules and FP modules. The time embedding and multi-scale latent representations are concatenated with the point features at the input of each corresponding SA and FP layer. We adopt the hybrid denoising score network parameterization following [3].
Results We validate our model’s results under two settings: normalizing data across the entire dataset and normalizing each shape individually. Samples from MLPCM are shown in Figure 3 and Figure  4, and quantitative results are provided in Table I and Table II. For airplanes, our results better preserve thin structures and global symmetry, with crisper wing contours and fewer broken tips. Competing methods often exhibit wing thickening, tail erosion, or rough leading edges, especially in PointFlow and ShapeGF, while DiT-3D improves the silhouette but still shows uneven point density around the tail. For chairs, our method maintains slender legs with consistent spacing and recovers backrest curvature and armrest topology with clear connectivity, while other methods produce floating clusters near the seat or excessively smoothed armrests. Although LION and PVD capture the overall shape, they lose fine perforations and edge sharpness. For cars, our generations show clearly defined wheel arch depressions and fewer surface pits, whereas other methods yield blotchy bodies with collapsed cavities or roof perforations. DPM and ShapeGF in particular display pronounced surface granularity and local self-intersections.
More visual results generated by our multi-scale latent consistency model are demonstrated in Figure 7. Overall, our model achieves sharper thin structures, cleaner topology, and more uniform point distributions, leading to higher structural fidelity than other methods.
@@
Note that TM stands for Teacher Model and LCM stands for Latent Consistency Models. MLPCM outperforms all baselines and achieves state-of-the-art performance across all categories and dataset versions. Compared to key baselines like DPM, PVD, and LION, our samples are diverse and have a better visual quality.
@@
TABLE I: Generation metrics (1-NNA↓) on airplane, chair, car categories from ShapeNet dataset from PointFlow [2]. Training and test data normalized globally into [-1, 1]. TM stands for Teacher Model and LCM stands for Latent Consistency Models.
@@
TABLE II: Generation results (1-NNA↓) on ShapeNet dataset from PointFlow [2]. All data normalized individually into [-1, 1].
@@
TABLE III: Results of joint train on 13 classes of ShapeNet-vol.
@@
TABLE IV: The sampling rates and generation quality of our model with various methods. Our approach achieves high-quality shape generation within 0.5 seconds while maintaining quality.
@@
TABLE V: Ablation studies on key 3D design components on the car category. PL and SL represent the point-level latents and shape-level latents, respectively, and 3DSA represents the 3D Spatial Attention Mechanism.
@@
TABLE VI: Ablation on the number of extra dimension $D_{l}$ of the latent points, on the car category.
@@
TABLE VII: Ablation study on the backbone of the encoder and decoder of hierarchical VAE and diffusion prior on the car category for Teacher Model. For DGCNN, we try different numbers of top-k nearest neighbors(knn).
[…]
DGCNN (knn=10) 67.66 57.08
DGCNN (knn=20) 68.54 57.60
PointTransformer 89.41 86.52
PointTransformerV2 80.09 79.57
PVCNN 51.17 48.92
TABLE VIII: Ablation study on 3D positional encoding on the car category for Teacher Model.
[…]
w/o 3D PE 51.60 49.38
with VPE 51.17 48.92
with cRPE 51.10 49.08
Figure 7: More visualization of Unconditional 3D shape generation via the proposed multi-scale latent consistency model on the airplane, car, and chair categories.
IV-D Multi-Class 3D Shape Generation
We validated the model’s generation performance on the Shapenet-Vol dataset, where each shape is normalized, meaning the point coordinates are bounded within [1,1][-1,1]. We report the quantitative results in Table III, comparing our method with various strong baselines under the same setting. We found that the proposed multi-scale latent diffusion model significantly outperforms all baselines, and MLPCM greatly improves sampling efficiency while maintaining performance.
IV-E Comparison with Nearest Neighbors and Voxel-guided Shape Synthesis
It is widely acknowledged that 1-NNA serves as the primary evaluation metric in this field, largely because no superior alternative currently exists. However, it can be misleading, as perfect memorization of the training set would yield an excellent score. Therefore, we adopt a more informative analysis that present the generated shape alongside its retrieved nearest neighbor, highlighting their structural and semantic similarity while also illustrating diversity in finer details. Specifically, Fig. 5 compares our generated shapes (top) with their retrieved Chamfer‐nearest neighbors (bottom). Global topology aligns closely, while fine details differ, confirming the diversity.
To supplement the conditional generation experiment to observe the generalization diversity of the model, We add voxel guided generation results as shown in Fig. 6. We begin by voxelizing the training set and fine-tuning our encoder networks to learn representations that can be accurately decoded back into the original shapes. For voxelized inputs processed with our point-cloud networks, we further sample surface points from the voxels to serve as input data.
IV-F Sampling Time
The proposed multi-scale latent diffusion model synthesizes shapes using 1000 steps of DDPM, while MLPCM synthesizes them using only 1-4 steps, generating high-quality shapes in under 0.5 seconds. This makes real-time interactive applications feasible. In Table IV, we compare the sampling time and quality of generating a point cloud sample (2048 points) from noise using different DDPM, DDIM [31] sampling methods, and MLPCM. When using $\leq 10$ steps, DDIM’s performance degrades significantly, whereas MLPCM can produce visually high-quality shapes in just a few steps or even one step.
IV-G Ablation Study
Ablation on the Key Design Components. We perform the ablation study on the Car category. For the configuration without multi-scale latent representations, we consider two scenarios: one using only point-level latent features to capture fine-grained details and another using only shape-level features to focus on global structure. This allows us to assess the impact of each representation scale individually. The model sizes are kept nearly identical across all settings to ensure fair comparison. As shown in Table V, the full model, integrating multi-scale latent representations and the 3D spatial attention mechanism, consistently outperforms the simplified settings across all evaluation metrics. This demonstrates the effectiveness of combining fine-grained local information with global shape context, along with the attention mechanism, in achieving more accurate and coherent reconstructions.
@@ IV-G Ablation Study
Ablation on Extra Dimension for Latent Points. We ablate the additional latent point dimension $D_{l}$ on the Car category in Table VI. We experimented with several different additional dimensions for the latent points, ranging from 0 to 5, where $D_{l}=1$ provided the best overall performance. As the additional dimensions increase, we observe a decrease in the 1-NNA score. We thus use $D_{l}=1$ for all other experiments.
Ablation on different backbones. We conduct an ablation study to evaluate different neural network architectures for processing point clouds in the encoder, decoder, and diffusion prior. The results, presented in Table VII, focus on the car category using the teacher model, consistent with other ablation studies. We test four popular backbone architectures: Point-Voxel CNN (PVCNN) [68], Dynamic Graph CNN (DGCNN) [21], PointTransformer [80], and PointTransformerV2 [81]. Among these, PVCNN demonstrates the best overall performance, making it our architecture of choice.
Ablation on 3D Positional Encoding. We conduct an ablation study to evaluate different 3D positional encoding methods. In the Voxel Positional Encoding (VPE) approach, we applied frequency-based sine-cosine 3D positional embeddings to the point-voxel branch in the SA module, following the design of DiT-3D [17]. For Contextual Relative Positional Encoding (cRPE), inspired by the Stratified Transformer [82], we enhanced the traditional positional encoding in the attention mechanism by adding a vector representing the positional offset between Query and Key during each attention computation. The experiments were performed on the car category, with results summarized in Table VIII. The findings show that omitting 3D positional encoding leads to significantly poorer generation results. While VPE and cRPE achieve comparable outcomes, we selected VPE as our 3D positional encoding method for its simplicity and effectiveness.
V Conclusion